The following information is provided about an open economy with a government. Use the Information to answer the questions that follow: C=450 +0.4Y 1 = 350 G= 150 X-70 2 = 35 +0.18 T 0.15Y Yf - 1550

Q.2.1 Calculate the level of autonomous spending in this economy. 2


1
Expert's answer
2020-12-03T05:20:54-0500

At equilibrium AS = AD

Y = C+I

Y = 450+0.4Y+350

Y = 450+350+0.4Y

Y-0.4Y = 800

0.6Y = 800

Y = 800/0.6

= 1333.33